DIFFERENTIAL ANALYSIS ky Manufacturing needs to obtain a gear-cutting machine, which can be purchased for Badons $75,000. Badonsky estimates that repair, maintenance, insurance, and property tax expense will be $20,000 for the machines 5-year life. At the end of the machines life, it will have no salvage value. As an alternative, Badonsky can lease the machine for 5 years for $18,000 per year. If the machine is leased, Badonsky is required to pay only for routine maintenance, which is estimated to be $8,000 over the machines life. All other costs will be paid by the lessor. Prepare a differential analysis to determine whether Badonsky should purchase or lease the machine.
Answer #1

Ans) Assuming the Rate of Return be default 12% in both cases;

In case of Purchasing-

Cost for purchasing- $75000

Maintenance Cost per year- $20,000/5 = $4000

Therefore, Present Value(PV) of Annuity for maintenance cost of $4000= $14,419

Formula; PV of Annuity= P{[1-(1+R)^-N] / r} = 4000{[1-(1+12)^-5] / 12} = $14,419

Therefore Total Amount = $75,000 + $14,419= $89,419

In case of Lease-

Cost per year= $18,000

Maintenance cost per year= $8000/5=$1600

Total cost = $18000+$1600= $19,600

PV of Annuity= P{[1-(1+R)^-N] / r} = 19600{[1-(1+12)^-5] / 12} = $70653.61

Conclusion- Lease will be a better option.